Saturday. And still a lot of things got done. – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

NEEM > Field diaries folder > UK diaries 2009 > Saturday. And still a ...

20 June 2009

Saturday. And still a lot of things got done.

The NEEM singers perform this summer's hit: The NEEM song.

The NEEM singers perform this summer's hit: The NEEM song.

All available hands were busy digging the cooling tunnel in the science trench. Some cut snow blocks, some hauled them out to the elevator in the drill trench, some took them out of the weatherport at the top of the elevator, and some drove the blocks away. A long chain that worked well.

The core is getting more brittle. Today one broke when it was pushed out of the core barrel. Still, the drillers and loggers were able to retrieve today’s production in a good way.

As blinds have been mounted in the main dome, we were able to dine in a pretend night. The darkness gave an altogether different feel to Saturday night. Going outside was a shocking experience. The light from the near midsummer Sun was blinding.

What we have done today:

  1. Drilling with the NEEM long drill.
  2. Logging. Last bag: 997. Depth: 548.35 m.
  3. No Processing. Processed depth: 329.45 m.
  4. Measurements in the CFA laboratory. Now 24 hours in two shifts.
  5. Grooming skiway taxi way and apron with tiller.
  6. Excavating cooling tunnel for science trench in front and to the left of core buffer. 6 m has been excavated, and we are half way.
  7. Adjusting saws at logging table. The circular saw was exchanged with a Makita band saw, and this gave good results. Loggers say that Makita is better than circular saw.
  8. Weighing and documenting pallets with retrograde cargo.
  9. Collecting samples at stable isotope pit.
  10. The mounting of blinds in the main dome is completed.
  11. Saturday night celebrated in darkness in candle light. We had an Asian dinner prepared by Nobohiko (Japan), JiWoong (Corea) and Cunde (China).

Ad.1: This is how the drillers report looks today:
"Stable drilling. Passed bag 1000! We drilled 17.49 m in 5 runs; drillers depth 542.18 m."

Ad.4: The CFA group is now working in two shifts, night and day. The measurements are running and so far they have measured from 7 to 33 m depth in the firn core. Melting of the firn became more difficult with depth, due to problems with air at the melt head. The team then jumped to 60m depth. Here melting and measurements went fine, and they are now slowly moving upwards to 50 m depth. The sulfate system is not working at the moment, but this is compensated by the U.K. fast I.C. which now includes sulfate. Some work has been done on tuning the water Picarro to the CFA.

Weather: In the morning broken cloud cover, in the afternoon and evening overcast and snow showers. -15°C to -9 °C, 15 knots from SW. Visibility: 5 miles, reduced to less than one mile during snow showers.

Field Leader, J.P. Steffensen

← Previous entry   Next entry →